OpenStudy (anonymous):

Which of the following |5-12i|? help please?

OpenStudy (anonymous):

\[|a+bi|=\sqrt{a^2+b^2}\]

OpenStudy (anonymous):

forget the \(i\) and forget the minus sign and compute \[\sqrt{5^2+12^2}\]